. From a box containing 4 black balls and 3 green balls, 4 balls are drawn in succession. Let Y be a random variable representing the number of green balls that occur.


1
Expert's answer
2022-02-28T13:22:12-0500

Let's denote B - black ball, G - green ball.

Sample space S is all possible outcomes.


"S=\{BBBB, BBBG, BBGB, BGBB, GBBB,"

"BBGG, BGBG, BGGB, GBGB, GBBG, GGBB,"

"BGGG, GBGG, GGBG, GGGB\}"

The possible values of the random variable "Y" are "0, 1, 2, 3."

We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2""\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & Z \\ \hline
 BBBB & 0 \\
 \hdashline
 BBBG & 1 \\
 \hdashline
 BBGB & 1 \\
 \hdashline
 BGBB & 1 \\
 \hdashline
 GBBB & 1 \\
 \hdashline
 BBGG & 2 \\
 \hdashline
 BGBG & 2 \\
 \hdashline
 BGGB & 2 \\
 \hdashline
 GBGB & 2 \\
 \hdashline
 GBBG & 2 \\
 \hdashline
 GGBB & 2 \\
 \hdashline
 BGGG & 1 \\
 \hdashline
 GBGG & 1 \\
 \hdashline
 GGBG & 1 \\
 \hdashline
 GGGB & 1 \\
 \hdashline
\end{array}"


Construct the probability distribution of the random variable


"P(BBBB)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{2}{5})(\dfrac{1}{4})=\dfrac{1}{35}"

"P(BBBG)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{2}{5})(\dfrac{3}{4})=\dfrac{3}{35}"

"P(BBGB)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{3}{5})(\dfrac{2}{4})=\dfrac{3}{35}"

"P(BGBB)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{3}{5})(\dfrac{2}{4})=\dfrac{3}{35}"


"P(GBBB)=\dfrac{3}{7}(\dfrac{4}{6})(\dfrac{3}{5})(\dfrac{2}{4})=\dfrac{3}{35}"

"P(BBGG)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{3}{5})(\dfrac{2}{4})=\dfrac{3}{35}"

"P(BGBG)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{3}{5})(\dfrac{2}{4})=\dfrac{3}{35}"

"P(BGGB)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{2}{5})(\dfrac{3}{4})=\dfrac{3}{35}"

"P(GBGB)=\dfrac{3}{7}(\dfrac{4}{6})(\dfrac{2}{5})(\dfrac{3}{4})=\dfrac{3}{35}"

"P(GBBG)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{3}{5})(\dfrac{2}{4})=\dfrac{3}{35}"


"P(GGBB)=\dfrac{3}{7}(\dfrac{2}{6})(\dfrac{4}{5})(\dfrac{3}{4})=\dfrac{3}{35}"

"P(BGGG)=\dfrac{4}{7}(\dfrac{3}{6})(\dfrac{2}{5})(\dfrac{1}{4})=\dfrac{1}{35}"

"P(GBGG)=\dfrac{3}{7}(\dfrac{4}{6})(\dfrac{2}{5})(\dfrac{1}{4})=\dfrac{1}{35}"

"P(GGBG)=\dfrac{3}{7}(\dfrac{2}{6})(\dfrac{4}{5})(\dfrac{1}{4})=\dfrac{1}{35}"

"P(GGGB)=\dfrac{3}{7}(\dfrac{2}{6})(\dfrac{1}{5})(\dfrac{4}{4})=\dfrac{1}{35}"


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 g & 0 & 1 & 2 & 3 \\ \hline
 p(g) & 1/35 & 12/35 & 18/35 & 4/35 
\end{array}"